x2Uxx-xyUxy+y2Uyy-xUx+yUy=8y/x (x≠0,y≠0) find the general solution.


1
Expert's answer
2021-02-16T05:55:45-0500

"\xi=\xi(x,y),\eta=\eta(x,y)"

"w(\xi,\eta)=u(x(\xi,\eta),y(\xi,\eta))\implies u(x,y)=w(\xi(x,y),\eta(x,y))"

"aw_{\xi\xi}+bw_{\xi\eta}+cw_{\eta\eta}=\varphi(\xi,\eta,w,w_{\xi},w_{\eta})"

"A=x^2,B=-xy,C=y^2"

"B^2-4AC=x^2y^2-4x^2y^2=-3x^2y^2<0"

This is an elliptic equation.

The roots of the characteristic polynomial:

"\lambda_1=\frac{B-i\sqrt{4AC-B^2}}{2A}=\frac{-xy-i\sqrt{4x^2y^2-x^2y^2}}{2x^2}=-\frac{y(1+i\sqrt{3})}{2x}"

"\lambda_2=\frac{B+i\sqrt{4AC-B^2}}{2A}=-\frac{y(1-i\sqrt{3})}{2x}"


"\frac{dy}{dx}=-\frac{y(1+i\sqrt{3})}{2x}\implies lny=-\frac{1+i\sqrt{3}}{2}lnx+lnc_1"


"c_1=yx^{(1+i\sqrt{3})/2}=\alpha"


"\frac{dy}{dx}=-\frac{y(1-i\sqrt{3})}{2x}"


"c_2=yx^{(1-i\sqrt{3})/2}=\beta"


"\xi=\frac{\alpha+\beta}{2}=y\sqrt{x}cos(\sqrt{3}lnx/2)"


"\eta=\frac{\alpha-\beta}{2i}=y\sqrt{x}sin(\sqrt{3}lnx/2)"


"\xi_x=\frac{y}{2\sqrt{x}}(cos(\sqrt{3}lnx/2)-\sqrt{3}sin(\sqrt{3}lnx/2))"


"\eta_x=\frac{y}{2\sqrt{x}}(sin(\sqrt{3}lnx/2)+\sqrt{3}cos(\sqrt{3}lnx/2))"


"\xi_y=\sqrt{x}cos(\sqrt{3}lnx/2)"


"\eta_y=\sqrt{x}sin(\sqrt{3}lnx/2)"


"\xi_{xx}=-\frac{y}{4x^{3/2}}(cos(\sqrt{3}lnx/2)-\sqrt{3}sin(\sqrt{3}lnx/2))+"

"+\frac{y\sqrt{3}}{4x^{3/2}}(-sin(\sqrt{3}lnx/2)-\sqrt{3}cos(\sqrt{3}lnx/2))="

"=\frac{y}{2x^{3/2}}(cos(\sqrt{3}lnx/2)+\sqrt{3}sin(\sqrt{3}lnx/2))"


"\eta_{xx}=\frac{y}{2x^{3/2}}(\sqrt{3}cos(\sqrt{3}lnx/2)-sin(\sqrt{3}lnx/2))"


"\xi_{yy}=\eta_{yy}=0"


"\xi_{xy}=\frac{1}{2\sqrt{x}}(cos(\sqrt{3}lnx/2)-\sqrt{3}sin(\sqrt{3}lnx/2))"


"\eta_{xy}=\frac{1}{2\sqrt{x}}(sin(\sqrt{3}lnx/2)+\sqrt{3}cos(\sqrt{3}lnx/2))"


Let: "k=cos(\sqrt{3}lnx/2), m=sin(\sqrt{3}lnx/2)"


"u_x=w_{\xi}\xi_x+w_{\eta}\eta_x"

"u_x=\frac{y}{2\sqrt{x}}(k-m\sqrt{3})w_{\xi}+\frac{y}{2\sqrt{x}}(m+k\sqrt{3})w_{\eta}"


"u_y=w_{\xi}\xi_y+w_{\eta}\eta_y"

"u_y=k\sqrt{x}w_{\xi}+m\sqrt{x}w_{\eta}"


"u_{xx}=w_{\xi\xi}\xi^2_x+2w_{\xi\eta}\xi_x\eta_x+w_{\eta\eta}\eta^2_x+w_{\xi}\xi_{xx}+w_{\eta}\eta_{xx}"

"u_{xx}=\frac{y^2}{4x}(k-m\sqrt{3})^2w_{\xi\xi}+\frac{y^2}{2x}(k-m\sqrt{3})(m+k\sqrt{3})w_{\xi\eta}+"

"+\frac{y^2}{4x}(m+k\sqrt{3})^2w_{\eta\eta}+\frac{y}{2x^{3/2}}(k+m\sqrt{3})w_{\xi}+\frac{y}{2x^{3/2}}(k\sqrt{3}-m)w_{\eta}"


"u_{yy}=w_{\xi\xi}\xi^2_y+2w_{\xi\eta}\xi_y\eta_y+w_{\eta\eta}\eta^2_y+w_{\xi}\xi_{yy}+w_{\eta}\eta_{yy}"

"u_{yy}=xk^2w_{\xi\xi}+2xkmw_{\xi\eta}+xm^2w_{\eta\eta}"


"u_{xy}=w_{\xi\xi}\xi_x\xi_y+w_{\xi\eta}(\xi_x\eta_y+\xi_y\eta_x)+w_{\eta\eta}\eta_x\eta_y+w_{\xi}\xi_{xy}+w_{\eta}\eta_{xy}"

"u_{xy}=\frac{y}{2}k(k-m\sqrt{3})w_{\xi\xi}+\frac{y}{2}(m(k-m\sqrt{3})+k(m+k\sqrt{3}))w_{\xi\eta}+"

"+\frac{y}{2}m(m+k\sqrt{3})w_{\eta\eta}+\frac{1}{2\sqrt{x}}(k-m\sqrt{3})w_{\xi}+\frac{1}{2\sqrt{x}}(m+k\sqrt{3})w_{\eta}"


Substituting the give values into the initial equation:


"\frac{3}{4}xy^2(w_{\xi\xi}+w_{\eta\eta})=\frac{8y}{x}-ym\sqrt{x}(w_{\xi}\sqrt{3}-w_{\eta})"


We also have:


"\xi^2+\eta^2=xy^2"


"x=e^{2tan^{-1}(\eta/\xi)/\sqrt{3}}"


"\frac{y}{x}=\frac{\sqrt{\xi^2+\eta^2}}{e^{tan^{-1}(\eta/\xi)/\sqrt{3}}}"


So, canonical form of equation:


"w_{\xi\xi}+w_{\eta\eta}=\frac{32}{3e^{tan^{-1}(\eta/\xi)/\sqrt{3}}\sqrt{\xi^2+\eta^2}}-\frac{4\eta}{3(\xi^2+\eta^2)}(w_{\xi}\sqrt{3}-w_{\eta})"